OpenStudy (anonymous):

4. If you drive 50 miles each week and your car gets 25 miles to the gallon with gas costing $3.00 a gallon, how much gas will you have to buy?

OpenStudy (unklerhaukus):

\[ 50 \frac{[\text{mile}]}{[\text{week}]} \div 25 \frac{\text{[mile]}}{[\text{gallon}]}\times{3}\frac{[$]}{[\text{gallon}]}\]

OpenStudy (anonymous):

umm... lets see

OpenStudy (anonymous):

75?

OpenStudy (anonymous):

nononononono

OpenStudy (anonymous):

125?

OpenStudy (unklerhaukus):

\[=50 \frac{[\text{mile}]}{[\text{week}]} \times \frac1{25} \frac{[\text{gallon}]}{\text{[mile]}}\times{3}\frac{[$]}{[\text{gallon}]}\]

OpenStudy (unklerhaukus):

\[=50\times\frac1{25}\times3 \frac{\cancel{[\text{mile}]}}{[\text{week}]} \times \frac{[\cancel{\text{gallon}]}}{\cancel{\text{[mile]}}}\times\frac{[$]}{\cancel{[\text{gallon}]}}\]

OpenStudy (anonymous):

Im not sure dude!